OpenStudy (anonymous):

Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make.

OpenStudy (anonymous):

@math92130

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

where did your work go it just disappered

OpenStudy (jack1):

2x + 4 = –3x + 14 2x + 4 +3x = –3x + 14 +3x 5x + 4 = + 14 5x + 4 -4 = + 14 -4 5x = + 10 x = 2

OpenStudy (jack1):

still works

OpenStudy (anonymous):

thank you